Hi The question is find the equations for exponential functions that pass through the given oair of points . round the coefficeints to 4 decimal places if necessary. ( -2,1) ( 1,5) f(x) =
Answer by stanbon(75887)

The question is find the equations for exponential functions that pass through the given pair of points . round the coefficeints to 4 decimal places if necessary. ( -2,1) ( 1,5) f(x) =
------------------------------------------------------
y = ab^x
5 = ab^1
1 = ab^-2
------------------
Divide the 1st equation by the 2nd to solve for "b":
5 = b^3
b = 5^(1/3)
------------
Substitute into 5 = ab^1 and solve for "a":
5 = a*5^(1/3)
a = 5^(2/3)
----------------------
Equation:
y = (5^(2/3))(5^(x/3)
===========================
Cheers,
stan H.